You are ordering softballs for 2 leagues. Pony league uses a ball that costs $2.75, junior league uses a ball that costs $3.25. Total balls bought is 80 for $245. How many for each league were bought?
Answer by MathPro(14)

Good evening,
Your problem:
You are ordering softballs for 2 leagues. Pony league uses a ball that costs $2.75, junior league uses a ball that costs $3.25. Total balls bought is 80 for $245. How many for each league were bought?
Call the number of balls that the pony league gets P and the number of balls that the junior league gets J. We are told that a total of 80 balls was bought. So this means that
(1) P+J=80
To digress, if you buy two items at $1.00 each the total cost is 2 x $1.00 = $2.00. Lets calculate the amount of money spent on each team. The price per ball is $2.75 for the ponies, so the total amount of money spent on this team is P x $2.75. The price per ball for the juniors is $3.25 so the total amount of money spent on this team is J x $3.25.
The total amount spent on both teams is the sum of the amount of money spent on each team:
Total amount= P x $2.75 + J x $3.25. We are given the total amount of money spent on both teams in the problem, $245, so we get
(2) 2.75P + 3.25J =245
Now we have two equations with two unknowns. Let’s solve these by substitution:
From equation (1) we get: P = 80 –J
Substitute for P in equation (2) : 2.75(80 –J) +J3.25 = 245
Simplify and solve: 220 +0.5J = 245, 0.5J = 25, J = 50
Substitute for J in equation (1) to get P + 50 = 80, P=30
Check the result by substituting for P and J in equation (2):
2.75(30) + 3.25(50)=245
The left-hand-side is equal to the right-hand-side, so the answer is correct.
Good Luck!
Selma from Rent-A-Whiz, an online tutoring and homework help website. Just visit www.rentawhiz.com to learn more. Currently, we are offering to solve one math problem for free. Just visit our website, select the homework tab, then click on the math problem submission link and quote coupon code 003 when you submit your problem.